Reports 1-1 of 1 Clear search Modify search
Injection system (EIB-SAS)
pillant - 9:12 Tuesday 22 January 2019 (44470) Print this report
Accelerometer 50Hz line

Since the EIB was suspended again, strong 50Hz and harmonic appeared in the strectrum of the accelerometer located on the bench.

Quadrants and photodiodes on the bench are not affected (attached plots).

Images attached to this report
Comments to this report:
fiori, derossi, paoletti - 12:41 Tuesday 19 February 2019 (44922) Print this report

After having investigated the accelerometer conditioning electronics in EERoom and found nothing, today we investigated the part on EIB.

We found the culprint was the low-noise coaxial 10-32 to BNC cable that happened to be damaged maybe during the last intervention.

Camilla replaced the cable, and the accelerometer now looks ok (Figure 1, purple is now).

Images attached to this comment
Search
×

Warning

×